OpenStudy (anonymous):

Solve the equation using the quadratic formula. 2x2 = -12x - 3. plz help

ganeshie8 (ganeshie8):

start by moving everything to left side

ganeshie8 (ganeshie8):

\(2x^2 = -12x -3\)

ganeshie8 (ganeshie8):

\(2x^2 + 12x +3 = 0\)

ganeshie8 (ganeshie8):

compare it wid standard quadratic equation : \(\large ax^2 + bx+c = 0\)

ganeshie8 (ganeshie8):

\(a = 2\) \(b = 12\) \(c = 3\)

ganeshie8 (ganeshie8):

plug them in quadratic formula : \(\large x = \frac{-b \pm \sqrt{b^2-4ac}}{2a}\)

ganeshie8 (ganeshie8):

\(\large x = \frac{-12 \pm \sqrt{12^2-4(2)(3)}}{2(2)}\)

ganeshie8 (ganeshie8):

simplify

OpenStudy (anonymous):

ok thanks

ganeshie8 (ganeshie8):

np :)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

2x2 = -12x - 3 2x2 = 4 -12 ÷ 4 = -3

OpenStudy (anonymous):

x=12(−6+30−−√),−12(6+30−−√)